jack is 6 years older than Jill. If four years ago Jack was 3 years less than twice Jill's age, how old is Jill now?
Answer by algebrahouse.com(1659)

x = Jill's age now
x + 6 = Jack's age now {Jack is 6 years older than Jill}

x - 4 = Jill's age 4 years ago
x + 2 = Jack's age 4 years ago {subtracted 4 from x + 6}

2(x - 4) - 3 = x + 2 {four years ago, Jack's age was equal to 3 years less than twice Jill's age}
2x - 8 - 3 = x + 2 {used distributive property}
2x - 11 = x + 2 {combined like terms}
x = 13 {subtracted x and added 11 to each side}

Jill is 13 now

For more help from me, visit:
www.algebrahouse.com